Find The sum of the first eight term of a linear sequence whose first term is 60 and whose last term is 46

Found 2 solutions by stanbon, greenestamps:
Answer by stanbon(75887)

Find The sum of the first eight term of a linear sequence whose first term is 60 and whose last term is 46
--------------------------
Find The sum of the first eight term of a linear sequence whose first term is 60 and whose last term is 46
-----
S(8) = (8/2)(60+46) = 4(106) = 424
-------------
Cheers
Stan H.
---------------

Answer by greenestamps(13200)



In a linear (arithmetic) sequence, the average of all the terms is the average of the first and last terms; in your problem the average is (60+46)/2 = 53.

Then since there are 8 terms, the sum of all the terms is 53*8 - = 424.